A business owner takes on a $10,000 loan to cover expenses. The interest rate on the loan is 6% per year. Each year, the business owner pays $1000 toward the loan. Find the balance on the loan after the second payment is made (round your answer to the nearest cent).
A) $9,176.00
B) $8,976.05
C) $6,788.08
D) $5,442.78

Answer by jim_thompson5910(35256)


Answer: Choice A) $9,176.00

-----------------------------------------------------------------------------
-----------------------------------------------------------------------------

Work Shown:

Year 1:
interest paid = (interest rate in decimal form)*(current balance)
interest paid = (0.06)*(10000)
interest paid = 600
-----------------
principal paid = (payment amount) - (interest paid)
principal paid = (1000) - (600)
principal paid = 400
-----------------
new balance = (old balance) - (principal paid)
new balance = (10000) - (400)
new balance = 9600
----------------------------------------------------------------------
Year 2:
interest paid = (interest rate in decimal form)*(current balance)
interest paid = (0.06)*(9600)
interest paid = 576
-----------------
principal paid = (payment amount) - (interest paid)
principal paid = (1000) - (576)
principal paid = 424
-----------------
new balance = (old balance) - (principal paid)
new balance = (9600) - (424)
new balance = 9176
----------------------------------------------------------------------
Table:
YearInterest PaidPrincipal PaidOld BalanceNew Balance
1$600$400$10,000$9,600
2$576$424$9,600$9,176